# Effective Rank Analysis and Regularization for Enhanced 3D Gaussian Splatting

### v2

@@ Abstract
3D reconstruction from multi-view images is one of the fundamental challenges in computer vision and graphics. Recently, 3D Gaussian Splatting (3DGS) has emerged as a promising technique capable of real-time rendering with high-quality 3D reconstruction. This method utilizes 3D Gaussian representation and tile-based splatting techniques, bypassing the expensive neural field querying. Despite its potential, 3DGS encounters challenges, including needle-like artifacts, suboptimal geometries, and inaccurate normals, due to the Gaussians converging into anisotropic Gaussians with one dominant variance. We propose using effective rank analysis to examine the shape statistics of 3D Gaussian primitives, and identify the Gaussians indeed converge into needle-like shapes with the effective rank 1. To address this, we introduce effective rank as a regularization, which constrains the structure of the Gaussians. Our new regularization method enhances normal and geometry reconstruction while reducing needle-like artifacts. The approach can be integrated as an add-on module to other 3DGS variants, improving their quality without compromising visual fidelity.
@@ 1 Introduction
However, while the flatness of Gaussians is necessary to make them align well with the surface, we argue that flatness alone is not sufficient for accurately representing surface geometry. Specifically, we observe that the majority of Gaussians converge into anisotropic forms with one dominant variance in 3DGS, effectively becoming needle-like with small scales along two of their axes. We identify this phenomenon as an important factor hindering accurate reconstruction, as needle-like Gaussians cover a negligible portion of the surface and create spiky artifacts. Disk-like Gaussians that cover non-negligible areas are actually needed for reconstructing the surface. However, previous methods do not properly distinguish between disk-like and needle-like Gaussians, as both have one of their scales near or exactly zero. In fact, we observe that in previous works, the majority of Gaussians converge into needle-like shapes.
To directly examine the shape statistics (whether their geometries are disk-like or needle-like) of 3D Gaussian primitives and understand their structural changes during training in a differentiable manner, we first propose performing effective rank analysis on the covariance matrices of Gaussians. The effective rank [25], which is a real-valued and differentiable extension of integer rank, can be utilized to monitor the training dynamics and structural transformations of Gaussian primitives. Indeed, our analysis reveals that the effective ranks of Gaussians approach an effective rank of 1 (erank-1), resulting in needle-like shapes in 3DGS and other methods, such as SuGaR [10] and 2DGS [13].
Additionally, we propose using effective rank as a regularization term to constrain the structure of the Gaussians. The differentiable nature of effective rank, with its concave logarithmic term providing stable gradients, makes it directly applicable to continuous optimization problems. Our new regularization method enhances normal and geometry reconstruction while reducing needle-like artifacts, particularly in novel view scenarios. Furthermore, our effective rank regularization can be applied as an add-on module to other 3DGS variants, improving their quality.
@@
  • We are firstly analyzing the dynamics of Gaussian primitive structures using effective rank in the optimizing process, discovering that Gaussians converge into anisotropic forms with one dominant variance.
@@
Figure 2: (green): Effective rank histograms for baseline methods 3DGS [16], SuGaR [10], and 2DGS [13], showing that Gaussian ranks are not optimally constrained for geometry reconstruction. (purple): The regularization term properly constrains the Gaussians, flattening them while preventing convergence into needle-like shapes.
@@ 2 Related work
Neural Radiance Fields (NeRF) [20] have revolutionized photo-realistic rendering from novel viewpoints by introducing a neural implicit representation of 3D scenes. This approach uses high-frequency positional encoding and differentiable volume rendering to achieve unprecedented realism. Enhancements to NeRF address challenges like anti-aliasing [1, 3], parameterizing unbounded scenes [2, 37], and training from in-the-wild images [19, 8, 29] through probabilistic transience modeling. Further improvements reduce training time and enhance rendering quality by incorporating low-rank tensor components [5].
Other research efforts have aimed for real-time rendering using alternative implicit models that do not rely on MLPs. Notable examples include sparse voxel grids [9] and multi-resolution hash encoding [21]. Despite these advancements, ray tracing methods are inherently slower than rasterization. To address this, 3D Gaussian Splatting (3DGS) [16] introduced a point-based rasterization technique for real-time, high-fidelity view synthesis. Inspired by EWA Volume Splatting [39], 3DGS uses a fully differentiable pipeline, representing 3D scenes with 3D Gaussians and performing volume splatting to known camera poses for rasterization.
@@
Recent advancements in 3D Gaussian Splatting (3DGS) have further propelled surface reconstruction. NeuSG [6] refines surface details using 3D Gaussian Splatting and neural implicit models. SuGaR [10] focuses on mesh extraction with SDF-based regularization and Poisson reconstruction. 2DGS [13] collapses 3D volumes into 2D Gaussian disks for view-consistent geometry and detailed mesh reconstruction. GaussianShader [15] enhances rendering quality in reflective surfaces using a shading function on 3D Gaussians. GOF [36] utilizes ray-Gaussian intersection for density estimation and geometric regularization. GIR [28] employs 3D Gaussians for inverse rendering, enabling accurate estimation of material properties, illumination, and geometry. These advancements showcase the potential of 3DGS for high-speed, detailed, and versatile surface reconstructions.
@@ 3 Preliminaries
3D Gaussian Splatting (3DGS) [16] represents a scene with a set of learnable 3D Gaussian primitives $\{\mathcal{G}_{k}\ |\ k=1,\cdots,K\}$, where each 3D Gaussian $\mathcal{G}_{k}$ consists of mean $\boldsymbol{\mu}_{k}\in\mathbb{R}^{3\times 1}$, covariance $\boldsymbol{\Sigma}_{k}\in\mathbb{R}^{3\times 3}$, point opacity $\alpha_{k}\in\left[0,1\right]$ and view-dependent color $c_{k}$ in spherical harmonics. Covariance matrix $\boldsymbol{\Sigma}_{k}=\mathbf{R}_{k}\mathbf{S}_{k}\mathbf{S}_{k}^{\top}\mathbf{R}_{k}^{\top}$ is positive semi-definite, where $\mathbf{S}_{k}=\mathrm{diag}(\mathbf{s}_{k})$ is a scaling matrix, $\mathbf{s}_{k}=(s_{k1};s_{k2};s_{k3})\in\mathbb{R}^{3\times 1}$ is a scale parameter, and $\mathbf{R}_{k}\in\mathbb{R}^{3\times 3}$ is a rotation matrix parameterized by a quaternion. A 3D Gaussian primitive can be represented in 3D space as:
@@ 3.2 Effective rank
The effective rank of the matrix $\mathbf{A}$ is concisely defined as $\text{erank}(\mathbf{A})=\text{exp}\{H(q_{1},q_{2},\cdots,q_{L})\},$ where $H(q_{1},q_{2},\cdots,q_{L})$ is the Shannon entropy given by
@@ 4 Method
In Section 4.1, we introduce effective rank analysis to inspect the geometries of Gaussians of 3DGS and its variants, shedding light on their underlying structures. Based on the findings from our effective rank analysis, we propose a novel effective rank regularization method in Section 4.2.
@@ 4.1 Effective rank analysis of 3D Gaussians
and the effective rank of a 3D Gaussian $\mathcal{G}_{k}$ with covariance matrix $\boldsymbol{\Sigma}_{k}$ is defined as follows:
@@
The effective rank, being a differentiable extension of integer rank, is a suitable tool for geometric analysis of 3D Gaussians since it jointly considers all of the scale parameters and can identify the relative scales of the three axes. The advantage of effective rank becomes more apparent when compared to recent works that only analyze individual or pair-wise variances of the 3D Gaussians [15]. Such approaches do not fully represent the geometry of Gaussians, potentially leading to planar and needle-like Gaussians being categorized together. For better understanding, we visualize effective ranks of sphere and 2D disks in Fig. 3.
With the distinct advantage of our approach, we can differentiate between needle-like Gaussians, which have effective ranks close to 1, and planar disk-like Gaussians. To reconstruct a scene with an accurate surface, we need Gaussians that represent a plane that aligns and concentrates well with the surface [30]. Ideally, 3D Gaussians with $\text{erank}(\mathcal{G}_{k})\approx 2$ are preferred, but Gaussians with effective rank smaller than 2 are also required for representing thin and elongated objects and patterns. Needle-like Gaussians with $\text{erank}(\mathcal{G}_{k})\approx 1$ are undesirable because they account for a negligible region of the surface and produce degenerate results in novel views.
The first row of Fig. 2 (green graph) shows the effective rank histogram for 3DGS during training. As the model converges, the number of 3D Gaussians with $\text{erank}(\mathcal{G}_{k})\approx 1$ increases, indicating overfitting without improvements in PSNR and Chamfer distance metrics (metrics are provided in the Appendix A.5, Table 5). This indicates that the majority of "flat" Gaussians (singular values close to 00) are actually needle-like ($\text{erank}(\mathcal{G}_{k})\approx 1$), rather than disk-like ($\text{erank}(\mathcal{G}_{k})\approx 2$). It is also interesting to note that 3DGS naturally forms a small mode at $\text{erank}(\mathcal{G}_{k})=2$, indicating a observed preference that can be further strengthened with our regularization.
@@ 4.2 Optimization
We adopt the revised version of the densification algorithm presented in [4, 36], which densifies Gaussians based on the summation of norms instead of the norm of the summation in Eq. 3 (further details in Appendix A.4). This change is particularly important for our regularization method. Unlike thin, needle-like Gaussians, disk-like Gaussians often do not meet the splitting criterion set by Eq. 3. This is because a disk-like Gaussian does not have a second axis with a much smaller variance than the axis with the largest variance. As a result, the gradient signals from each pixel are generally smaller compared to those from needle-like Gaussians. Furthermore, since disk-like Gaussians typically cover more pixel space, unaligned signals tend to cancel each other out. In contrast, the revised densification algorithm facilitates the splitting of Gaussians with our regularization. However, note that due to the efficacy of disk-like Gaussians in reconstructing the surface compared to needle-like ones, our method still requires about 10% fewer Gaussians than the baseline [16].
@@ 5.1 Implementation
The regularization hyperparameter $\lambda_{\text{erank}}=0.01$ is used for all training. For other components belonging to the baselines, we use the same settings as described in the corresponding papers. All experiments are conducted on a Tesla V100 GPU. For mesh extraction, truncated signed distance function (TSDF) fusion with Open3D [38] is used, with details in the Appendix A.3.
@@ 5.2 Comparison
Our method is applicable to other baselines as an add-on term. Therefore, we compare baselines with and without our regularization. We choose SuGaR, 2DGS, and GOF as our baselines, works that focus on better geometry reconstruction, along with the original 3D Gaussian Splatting. All of the experiments are performed with the proposed setting of the original paper.
@@
Table 1 presents the quantitative results of geometry reconstruction on the DTU dataset. We report the Chamfer distance for each scene, along with the mean Chamfer distance and mean PSNR. The “+e” symbol indicates the addition of effective rank regularization (with fixed densification) to the baseline methods.
@@
Figure 4 shows mesh reconstruction results, where baseline methods often exhibit empty holes in the reconstructed meshes. Our regularization term enforces disk-like Gaussians, reducing such holes and proving advantageous for surface reconstruction.
Figure 5 and the first row of Figure 1 display normal reconstruction results. In Fig. 5, the resulting image from GOF shows spiky artifacts and a hollow surface on the pear. Similarly to the mesh results, needle-like Gaussians often fail to cover the entire area, leaving empty holes or transparent regions, resulting in hollow or incomplete reconstructions. The effective rank regularization mitigates these noisy artifacts, leading to a more accurate reconstruction of the underlying geometry.
@@
As shown in Fig. 7 and Table 4, efficacy of disk-like Gaussians in 3D reconstruction, compared to needle-like Gaussians, leads to a better memory footprint. The average storage usage for the DTU and Mip-NeRF360 datasets is reported in Table 4.
@@ 5.3 Ablations
Our method comprises two key components: (a) the fixed densification (ADC) algorithm and (b) effective rank regularization. We performed an ablation study on these components to observe their performance gains compared to the naive 3DGS method. Table 2 shows the Chamfer distance and PSNR measured on the DTU dataset. The results indicate that both components contribute to performance gains in geometry reconstruction and novel view synthesis tasks. Additionally, incorporating techniques such as depth distortion loss [13, 36] can further enhance the best-performing model (row +a+b+c). These techniques are discussed in Appendix A.2.
@@ 6 Conclusion
Junha Hyung and Susung Hong conducted this work during the internship at NAVER AI Lab. The NAVER Smart Machine Learning (NSML) platform [17] had been used for experiments.
@@ A.1 Broader impact
The broader impact of our work on 3D reconstruction lies in its potential to advance various fields such as virtual and augmented reality, medical imaging, and digital content creation by enabling more efficient and high-quality 3D model generation. However, like any powerful technology, it also presents potential risks and avenues for misuse. For instance, enhanced 3D reconstruction techniques could be exploited to create deepfakes or unauthorized reproductions of proprietary designs, posing ethical and legal challenges. To mitigate these risks, we propose implementing strict usage guidelines to ensure the integrity and rightful use of 3D models. We aim to maximize the positive impact of our research while minimizing potential negative consequences.

### v3

@@ Abstract
3D reconstruction from multi-view images is one of the fundamental challenges in computer vision and graphics. Recently, 3D Gaussian Splatting (3DGS) has emerged as a promising technique capable of real-time rendering with high-quality 3D reconstruction. This method utilizes 3D Gaussian representation and tile-based splatting techniques, bypassing the expensive neural field querying. Despite its potential, 3DGS encounters challenges such as needle-like artifacts, suboptimal geometries, and inaccurate normals caused by the Gaussians converging into anisotropic shapes with one dominant variance. We propose using the effective rank analysis to examine the shape statistics of 3D Gaussian primitives, and identify the Gaussians indeed converge into needle-like shapes with the effective rank 1. To address this, we introduce the effective rank as a regularization, which constrains the structure of the Gaussians. Our new regularization method enhances normal and geometry reconstruction while reducing needle-like artifacts. The approach can be integrated as an add-on module to other 3DGS variants, improving their quality without compromising visual fidelity. The project page is available at https://junhahyung.github.io/erankgs.github.io/.
@@ 1 Introduction
While the flatness of Gaussians is necessary for proper alignment with the surface, we argue that flatness alone is insufficient for accurately representing surface geometry. Gaussians should also avoid being needle-like or highly anisotropic, where one variance dominates the others. Needle-like Gaussians hinder accurate reconstruction as they cover only a negligible portion of the surface and produce spiky artifacts. Instead, disk-like Gaussians are preferable as they cover meaningful areas and contribute effectively to surface reconstruction.
However, existing methods fail to adequately differentiate between disk-like and needle-like Gaussians, as both can exhibit one scale that is near or exactly zero. Empirically, we find that most Gaussians converge into anisotropic forms, becoming needle-like with small scales along two axes due to the absence of appropriate regularization mechanisms.
To directly examine the shape statistics (whether their geometries are disk-like or needle-like) of 3D Gaussian primitives and understand their structural changes during training, we first propose performing the effective rank analysis on the covariance matrices of Gaussians. The effective rank [25], which is a real-valued and differentiable extension of the integer rank, can be used to monitor the training dynamics and structural transformations of Gaussian primitives. Indeed, our analysis reveals that the effective ranks of Gaussians approach an effective rank of 1 (erank-1), resulting in needle-like shapes in 3DGS and other methods, such as SuGaR [10] and 2DGS [13].
Additionally, we propose using the effective rank as a regularization term to constrain the structure of the Gaussians. The differentiable nature of effective rank, with its concave logarithmic term providing stable gradients, makes it directly applicable to continuous optimization problems. Our new regularization method enhances normal and geometry reconstruction while reducing needle-like artifacts, particularly in novel view scenarios. Furthermore, our effective rank regularization can be applied as an add-on module to other 3DGS variants, improving their quality.
@@
  • We are firstly analyzing the dynamics of Gaussian primitive structures using the effective rank in the optimizing process, discovering that Gaussians converge into anisotropic forms with one dominant variance.
@@
Figure 2: (Green): The effective rank histograms for baseline methods 3DGS [16], SuGaR [10], and 2DGS [13], showing that Gaussian ranks are not optimally constrained for geometry reconstruction. (Purple): The regularization term properly constrains the Gaussians, flattening them while preventing convergence into needle-like shapes.
@@ 2 Related work
NeRFs [20] have revolutionized photo-realistic rendering from novel viewpoints by introducing a neural implicit representation of 3D scenes. This approach uses high-frequency positional encoding and differentiable volume rendering to achieve unprecedented realism. Enhancements to NeRF address challenges like anti-aliasing [1, 3], parameterizing unbounded scenes [2, 37], and training from in-the-wild images [19, 8, 29] through probabilistic transience modeling. Further improvements reduce training time and enhance rendering quality by incorporating low-rank tensor components [5].
Other research efforts have aimed for real-time rendering using alternative implicit models that do not rely on MLPs. Notable examples include sparse voxel grids [9] and multi-resolution hash encoding [21]. Despite these advancements, ray tracing methods are inherently slower than rasterization. To address this, 3DGS [16] introduced a point-based rasterization technique for real-time, high-fidelity view synthesis. Inspired by EWA Volume Splatting [39], 3DGS uses a fully differentiable pipeline, representing 3D scenes with 3D Gaussians and performing volume splatting to known camera poses for rasterization.
@@
3D Gaussian Splatting
Recent advancements in 3DGS have further propelled surface reconstruction. NeuSG [6] refines surface details using 3DGS and neural implicit models. SuGaR [10] focuses on mesh extraction with SDF-based regularization and Poisson reconstruction. 2DGS [13] collapses 3D volumes into 2D Gaussian disks for view-consistent geometry and detailed mesh reconstruction. GaussianShader [15] enhances rendering quality in reflective surfaces using a shading function on 3D Gaussians. GOF [36] utilizes ray-Gaussian intersection for density estimation and geometric regularization. GIR [28] employs 3D Gaussians for inverse rendering, enabling accurate estimation of material properties, illumination, and geometry. These advancements showcase the potential of 3DGS for high-speed, detailed, and versatile surface reconstructions.
@@ 3 Preliminaries
3DGS [16] represents a scene with a set of learnable 3D Gaussian primitives $\{\mathcal{G}_{k}\ |\ k=1,\cdots,K\}$, where each 3D Gaussian $\mathcal{G}_{k}$ consists of mean $\boldsymbol{\mu}_{k}\in\mathbb{R}^{3\times 1}$, covariance $\boldsymbol{\Sigma}_{k}\in\mathbb{R}^{3\times 3}$, point opacity $\alpha_{k}\in\left[0,1\right]$ and view-dependent color $c_{k}$ in spherical harmonics. Covariance matrix $\boldsymbol{\Sigma}_{k}=\mathbf{R}_{k}\mathbf{S}_{k}\mathbf{S}_{k}^{\top}\mathbf{R}_{k}^{\top}$ is positive semi-definite, where $\mathbf{S}_{k}=\mathrm{diag}(\mathbf{s}_{k})$ is a scaling matrix, $\mathbf{s}_{k}=(s_{k1};s_{k2};s_{k3})\in\mathbb{R}^{3\times 1}$ is a scale parameter, and $\mathbf{R}_{k}\in\mathbb{R}^{3\times 3}$ is a rotation matrix parameterized by a quaternion. A 3D Gaussian primitive can be represented in 3D space as:
@@ 3.2 Effective rank
The effective rank [25] of the matrix $\mathbf{A}$ is concisely defined as $\text{erank}(\mathbf{A})=\text{exp}\{H(q_{1},q_{2},\cdots,q_{L})\},$ where $H(q_{1},q_{2},\cdots,q_{L})$ is the Shannon entropy given by
@@ 4 Method
In Section 4.1, we introduce the effective rank analysis to inspect the geometries of Gaussians of 3DGS and its variants, shedding light on their underlying structures. Based on the findings from our effective rank analysis, we propose a novel effective rank regularization method in Section 4.2.
@@ 4.1 Effective rank analysis of 3D Gaussians
and the effective rank of a 3D Gaussian $\mathcal{G}_{k}$ with the covariance matrix $\boldsymbol{\Sigma}_{k}$ is defined as follows:
@@
The effective rank, being a differentiable extension of an integer rank, is a suitable tool for geometric analysis of 3D Gaussians since it jointly considers all of the scale parameters and can identify the relative scales of the three axes. The advantage of effective rank becomes more apparent when compared to recent works that only analyze individual or pair-wise variances of the 3D Gaussians [15]. Such approaches do not fully represent the geometry of Gaussians, potentially leading to planar and needle-like Gaussians being categorized together. For better understanding, we visualize the effective ranks of a sphere and 2D disks in Fig. 3.
With the distinct advantage of our approach, we can differentiate between needle-like Gaussians, which have effective ranks close to 1, and planar disk-like Gaussians. To reconstruct a scene with an accurate surface, we need Gaussians that represent a plane that aligns and concentrates well with the surface [30]. Ideally, 3D Gaussians with $\text{erank}(\mathcal{G}_{k})\approx 2$ are preferred, but Gaussians with an effective rank smaller than 2 are also required for representing thin and elongated objects and patterns. However, the needle-like Gaussians with $\text{erank}(\mathcal{G}_{k})\approx 1$ are undesirable because they account for a negligible region of the surface and produce degenerate results in novel views.
The first row of Fig. 2 (green graph) shows the effective rank histogram for 3DGS during training. As the model converges, the number of 3D Gaussians with $\text{erank}(\mathcal{G}_{k})\approx 1$ increases, indicating overfitting without improvements in PSNR and Chamfer distance metrics (metrics are provided in the Appendix A.5, Table 5). This indicates that the majority of "flat" Gaussians (singular values close to 00) are actually needle-like ($\text{erank}(\mathcal{G}_{k})\approx 1$), rather than disk-like ($\text{erank}(\mathcal{G}_{k})\approx 2$). It is also interesting to note that 3DGS naturally forms a small mode at $\text{erank}(\mathcal{G}_{k})=2$, indicating an observed preference that can be further strengthened with our regularization.
@@ 4.2 Optimization
We adopt the revised version of the densification algorithm presented in [4, 36], which densifies Gaussians based on the summation of norms instead of the norm of the summation in Eq. 3 (further details in Appendix A.4). This change is particularly important for our regularization method. Disk-like Gaussians, unlike needle-like ones, often fail to satisfy the splitting criterion set by Eq. 3 because their axes with larger variances produce smaller gradient signals per pixel. Moreover, since disk-like Gaussians typically cover more pixel space, unaligned signals tend to cancel each other out. In contrast, the revised densification algorithm facilitates the splitting of disk-like Gaussians. Notably, due to their superior ability to reconstruct surfaces compared to needle-like Gaussians, our method requires approximately 10% fewer Gaussians than the baseline [16].
@@ 5.1 Implementation
The regularization hyperparameter $\lambda_{\text{erank}}=0.01$ is used for all training. For other components belonging to the baselines, we use the same settings as described in the corresponding papers. All experiments are conducted on a Tesla V100 GPU. For mesh extraction, the truncated signed distance function (TSDF) fusion with Open3D [38] is used, with details in the Appendix A.3.
@@ 5.2 Comparison
Our method is applicable to other baselines as an add-on term. Therefore, we compare baselines with and without our regularization. We choose SuGaR, 2DGS, and GOF as our baselines, works that focus on better geometry reconstruction, along with the original 3DGS. All of the experiments are performed with the proposed setting of the original paper.
@@
Table 1 presents the quantitative results of geometry reconstruction on the DTU dataset. We report the Chamfer distance for each scene, along with the mean Chamfer distance and mean PSNR. The “+e” symbol indicates the addition of the effective rank regularization (with fixed densification) to the baseline methods.
@@
Fig. 4 shows mesh reconstruction results, where baseline methods often exhibit empty holes in the reconstructed meshes. Our regularization term enforces disk-like Gaussians, reducing such holes and proving advantageous for surface reconstruction.
Fig. 5 and the first row of Fig. 1 display normal reconstruction results. In Fig. 5, the resulting image from GOF shows spiky artifacts and a hollow surface on the pear. Similarly to the mesh results, needle-like Gaussians often fail to cover the entire area, leaving empty holes or transparent regions, resulting in hollow or incomplete reconstructions. The effective rank regularization mitigates these noisy artifacts, leading to a more accurate reconstruction of the underlying geometry.
@@
As shown in Fig. 7 and Table 4, the efficacy of disk-like Gaussians in 3D reconstruction, compared to needle-like Gaussians, leads to a better memory footprint. The average storage usage for the DTU and Mip-NeRF360 datasets is reported in Table 4 (Appendix A.5).
@@ 5.3 Ablations
Our method comprises two key components: (a) the fixed densification (ADC) algorithm and (b) the effective rank regularization. We performed an ablation study on these components to observe their performance gains compared to the naive 3DGS method. Table 2 shows the Chamfer distance and PSNR measured on the DTU dataset. The results indicate that both components contribute to performance gains in geometry reconstruction and novel view synthesis tasks. Additionally, incorporating techniques such as depth distortion loss [13, 36] can further enhance the best-performing model (row +a+b+c). These techniques are discussed in the Appendix A.2.
@@ 6 Conclusion
Junha Hyung and Susung Hong conducted this work during the internship at NAVER AI Lab. The NAVER Smart Machine Learning (NSML) platform [17] had been used for experiments. This work was supported by KAIST-NAVER hypercreative AI center. This work was partly supported by Institute for Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program (KAIST)). This work was also partly supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1A2B5B02001913).
@@ A.1 Broader impact
The broader impact of our work on 3D reconstruction lies in its potential to advance various fields such as virtual and augmented reality, medical imaging, and digital content creation by enabling more efficient and high-quality 3D model generation. However, as with any advanced technology, it also presents potential risks and avenues for misuse. For instance, enhanced 3D reconstruction techniques could be exploited to create deepfakes or unauthorized reproductions of proprietary designs, posing ethical and legal challenges. To mitigate these risks, we propose implementing strict usage guidelines to ensure the integrity and rightful use of 3D models. We aim to maximize the positive impact of our research while minimizing potential negative consequences.